# Concentration and fluctuations of sine–Gordon measure around topological multi-soliton manifold

### v1

@@ 7.1. Partition into a typical event and large deviation events
Let $Q\in\mathbb{Z}$ with $Q>0$, $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$, $d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$, as in (6.28) and (6.29), and let $F$ be a bounded function. Then
$\int F(\phi)\rho_{\varepsilon}^{Q}(d\phi)=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)+\textup{O}(e^{-c\log\frac{1}{\varepsilon}}).$
@@
Note that, using (6.28), we have
$\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon})<\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)+\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon})\geq\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)$
$=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon})<\delta_{\varepsilon}\}\cap\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\geq d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)$
$\hphantom{X}+\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon})<\delta_{\varepsilon}\}\cap\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\geq d_{\varepsilon}})\geq\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)+\textup{O}(e^{-c\log\frac{1}{\varepsilon}})$
$=\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}+\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2}+\textup{O}(e^{-c\log\frac{1}{\varepsilon}}).$ (7.1)
Since $\mathcal{M}_{Q}^{\varepsilon,\geq d_{\varepsilon}}\subset\mathcal{M}_{Q}^{\varepsilon}$
$\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\geq d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi).$ (7.2)
Using $\mathcal{M}_{Q}^{\varepsilon}=\mathcal{M}_{Q}^{\varepsilon,\geq d_{\varepsilon}}\cup\mathcal{M}_{Q}^{\varepsilon,<d_{\varepsilon}}$, we have
$\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon})<\delta_{\varepsilon}\}\cap\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\geq d_{\varepsilon}})\geq\delta_{\varepsilon}\}\subset\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,<d_{\varepsilon}})<\delta_{\varepsilon}\}.$
This implies, together with (6.29), that
$\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2}\lesssim\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,<d_{\varepsilon}})<\delta_{\varepsilon}\}\big)\lesssim e^{-c\cdot\varepsilon^{-\frac{1}{2}+}}.$ (7.3)
Combining (7.1), (7.2), and (7.3) yields
(7.1) $=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\geq d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)+\textup{O}(e^{-c\log\frac{1}{\varepsilon}})+\textup{O}(e^{-c\cdot\varepsilon^{-\frac{1}{2}+}})$
$=\mathcal{J}_{1}+\textup{O}(e^{-c\log\frac{1}{\varepsilon}}).$ (7.4)
We further decompose the leading-order term $\mathcal{J}_{1}$ as follows. Using (6.30), we have
$\mathcal{J}_{1}=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)+\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq d_{\varepsilon}})<\delta_{\varepsilon}\}\setminus\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)$ (7.5)
From the tail bound (6.32), we obtain
$\mathcal{J}_{1}=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)+\textup{O}(e^{-c\cdot\varepsilon^{-\frac{1}{2}+}}).$ (7.6)
Combining (7.1), (7.4), and (7.6) completes the proof.
Remark 7.2.
The large deviation estimates (6.28) and (6.29) imply that most of the probability mass is concentrated on the set $\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\geq d_{\varepsilon}})<\delta_{\varepsilon}\}$ in (7.2). We note, however, that this set is too large to prove the central limit theorem, since the error term is too large to control on this set. In particular, in (7.13) the first-order term in the Taylor expansion blows up on this set, with size $O(\varepsilon^{0-})$. Therefore, we need to perform a more careful decomposition in (7.5) in order to carry out the Taylor expansion and control the error on the slightly smaller set $\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}$ in (7.6).
7.2. Coordinate changes
@@
As shown in the previous subsection, we prove that as $\varepsilon\to 0$, most of the probability mass concentrates in the region,
$\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\},$
@@
where the field $\phi$ is close to the manifold and the solitons are well separated, that is, there is essentially no collision as $d_{\varepsilon}\to\infty$. See the definition of $\mathcal{M}^{\varepsilon,\geq d}_{Q}$ in (4.10).
To prove the central limit theorem (Theorem 1.4), we first perform the following change of coordinates using the disintegration formula in Lemma 4.3.
Remark 7.3.
As discussed in Remark 4.4, the disintegration formula in tangential and normal coordinates has been widely used in settings where the energy functional admits explicit minimizers. However, we emphasize again that, to the best of our knowledge, this is the first time the disintegration formula is applied in a regime where no minimizers exist, and the framework is set up around a multi-soliton manifold.
@@
Let $Q\in\mathbb{Z}$ with $Q>0$, $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$, $d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$, as in (6.28) and (6.29), and let $F$ be a bounded, continuous function. Then
$\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\sqrt{\varepsilon}^{-1}(\phi-\pi^{\varepsilon}(\phi)))\rho_{\varepsilon}^{Q}(d\phi)$
$=\overline{Z}_{\varepsilon}^{-1}\int\cdots\int_{U_{\varepsilon}}F(v)e^{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1}\,\dots,\xi_{Q}}(\sqrt{\varepsilon}v)}\textup{Det}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)\,\nu^{\perp}_{\xi_{1},\dots,\xi_{Q}}(dv)\,d\xi_{1}\dots,d\xi_{Q},$ (7.7)
where $\pi^{\varepsilon}$ is the projection onto $\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}}$ , defined in (4.15), $\nu^{\perp}_{\xi_{1},\dots,\xi_{Q}}$ is the Gaussian measure in Lemma 5.4, and
$U_{\varepsilon}:=\big\{(\xi_{1},\dots,\xi_{Q},v)\in\Delta_{Q}\times V_{\xi_{1},\dots,\xi_{Q}}:\|\sqrt{\varepsilon}v\|_{L^{2}}<\delta_{\varepsilon}\;\textup{and}\;\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\big\}$
@@
Here $\Delta_{Q}=\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}$. In (7.7), the partition function $\overline{Z}_{\varepsilon}$ represents the integral $\int\cdots\int$ over $U_{\varepsilon}$ without $F$, and $W_{\xi_{1},\dots,\xi_{Q},\sqrt{\varepsilon}v}$ is the Weingarten map, defined in (4.20).
@@
Using the disintegration formula in Lemma 4.3 with the coordinate $\phi=m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}+\sqrt{\varepsilon}v$ in (4.14), we decompose the integral into its tangential and normal components as follows
$Z[F]:$ $=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\sqrt{\varepsilon}^{-1}(\phi-\pi^{\varepsilon}(\phi)))\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\mu_{\varepsilon}^{Q}(d\phi)$
$=\int\cdots\int_{U_{\varepsilon}}F(v)\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}1-\cos(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}+\sqrt{\varepsilon}v)dx\Big\}$
@@
$\hskip 76.82243pt\cdot\exp\Big\{-\frac{1}{2\varepsilon}\|\partial_{x}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}\|_{L^{2}}^{2}-\frac{1}{\varepsilon}\langle(-\partial_{x}^{2})m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},\sqrt{\varepsilon}v\rangle_{L^{2}}\Big\}$
@@
$\hskip 76.82243pt\cdot\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)\,\mu^{\perp}_{\xi_{1},\dots,\xi_{Q}}(dv)\,d\sigma(\xi_{1},\dots,\xi_{Q}),$ (7.8)
@@
where $\pi^{\varepsilon}(\phi)=m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}$ with $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}$, and $\mu^{\perp}_{\xi_{1},\dots,\xi_{k}}$ is the Brownian bridge with covariance $(-\partial_{x}^{2})^{-1}$ from 0 to 0, on the normal space, and $d\sigma$ is the surface measure defined in (4.19).
Under the coordinate $\phi=m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}+\sqrt{\varepsilon}v$ in (4.14), we perform a Taylor expansion
$\int_{-L_{\varepsilon}}^{L_{\varepsilon}}1-\cos\phi\,dx$ $=\int_{-L_{\varepsilon}}^{L_{\varepsilon}}1-\cos(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})dx+\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\sin(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})\sqrt{\varepsilon}vdx$
$\hphantom{X}+\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\cos(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})(\sqrt{\varepsilon}v)^{2}dx+\mathcal{E}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v),$ (7.9)
where $\mathcal{E}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)=\int_{-L_{\varepsilon}}^{L_{\varepsilon}}O\big((\sqrt{\varepsilon}v)^{3}\big)dx$. By combining the Cameron–Martin term
$\frac{1}{2}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}\phi|^{2}dx=\frac{1}{2}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}|^{2}dx+\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(-\partial_{x}^{2}m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})\sqrt{\varepsilon}vdx+\frac{1}{2}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}(\sqrt{\varepsilon}v)|^{2}dx$
with the interaction terms (7.9), we obtain
$-\frac{1}{\varepsilon}E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}+\sqrt{\varepsilon}v)$ $=-\frac{1}{\varepsilon}E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})-\frac{1}{\varepsilon}\langle\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}),\sqrt{\varepsilon}v\rangle$
@@
$\hphantom{X}-\frac{1}{2\varepsilon}\langle\sqrt{\varepsilon}v,(-\partial_{x}^{2}+\cos m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})\sqrt{\varepsilon}v\rangle$
@@
$\hphantom{X}+\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v),$ (7.10)

### v2

@@ 7. Change of coordinates
Let $Q\in\mathbb{Z}$ with $Q>0$, $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$, $d_{\varepsilon}=\big|\log(\varepsilon\log\frac{1}{\varepsilon})\big|$, as in (6.29) and (6.30), and let $F$ be a bounded function. Then
$\int F(\phi)\rho_{\varepsilon}^{Q}(d\phi)=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)+\textup{O}(e^{-c\log\frac{1}{\varepsilon}}).$
@@
Write $A:=\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon})<\delta_{\varepsilon}\}$ and $B:=\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq d_{\varepsilon}})<\delta_{\varepsilon}\}$.
Since $\mathcal{M}_{Q}^{\varepsilon}=\mathcal{M}_{Q}^{\varepsilon,\geq d_{\varepsilon}}\cup\mathcal{M}_{Q}^{\varepsilon,<d_{\varepsilon}}$,
$A\cap B=B,\qquad A\cap B^{c}\subset\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,<d_{\varepsilon}})<\delta_{\varepsilon}\}.$
Decomposing the full set as $(A\cap B)\cup(A\cap B^{c})\cup A^{c}$, then applying (6.29) to $A^{c}$, and (6.30) to $A\cap B^{c}$, we obtain the lemma.
@@
The above lemma shows that as $\varepsilon\to 0$, most of the probability mass concentrates in the region,
$\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq d_{\varepsilon}})<\delta_{\varepsilon}\},$
@@
where the field $\phi$ is close to the manifold and the solitons are well separated, that is, there is essentially no collision as $d_{\varepsilon}\to\infty$.
To prove the central limit theorem (Theorem 1.4), we first perform the following change of coordinates using the disintegration formula in Lemma 4.2. As discussed in Remark 4.3, this is the first time the disintegration formula is applied in a regime with no minimizer, and in a setting around a multi-soliton manifold. Recall that $\pi^{\varepsilon}$ denotes the projection onto $\mathcal{M}_{Q}^{\varepsilon,\,\geq d_{\varepsilon}}$ defined in (4.15), and we write $\Delta_{Q}=\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}$.
@@
Let $Q\in\mathbb{Z}$ with $Q>0$, $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$, $d_{\varepsilon}=\big|\log(\varepsilon\log\frac{1}{\varepsilon})\big|$, as in (6.29) and (6.30), and let $F$ be a bounded, continuous function. Then
$\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\sqrt{\varepsilon}^{-1}(\phi-\pi^{\varepsilon}(\phi)))\rho_{\varepsilon}^{Q}(d\phi)$
(7.1) $=\overline{Z}_{\varepsilon}^{-1}\int\cdots\int_{U_{\varepsilon}}F(v)e^{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1}\,\dots,\xi_{Q}}(\sqrt{\varepsilon}v)}\textup{Det}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)\,\nu^{\perp}_{\xi_{1},\dots,\xi_{Q}}(dv)\,d\xi_{1}\dots,d\xi_{Q},$
where $\nu^{\perp}_{\xi_{1},\dots,\xi_{Q}}$ is the Gaussian measure in Lemma 5.4, and
$U_{\varepsilon}:=\big\{(\xi_{1},\dots,\xi_{Q},v)\in\Delta_{Q}\times V_{\xi_{1},\dots,\xi_{Q}}:\|\sqrt{\varepsilon}v\|_{L^{2}}<\delta_{\varepsilon}\;\textup{and}\;\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\big\}$
@@
In (7.1), the partition function $\overline{Z}_{\varepsilon}$ represents the integral over $U_{\varepsilon}$ with $F=1$, and $W_{\xi_{1},\dots,\xi_{Q},\sqrt{\varepsilon}v}$ is the Weingarten map, defined in (4.21).
@@
Using the disintegration formula in Lemma 4.2 with the coordinate
$\phi=\pi^{\varepsilon}(\phi)+\sqrt{\varepsilon}v,\qquad\pi^{\varepsilon}(\phi)=m^{\varepsilon}_{*}:=m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}$
we decompose the integral into its tangential and normal components as follows
$Z[F]:$ $=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\sqrt{\varepsilon}^{-1}(\phi-\pi^{\varepsilon}(\phi)))\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\mu_{\varepsilon}^{Q}(d\phi)$
$=\int\cdots\int_{U_{\varepsilon}}F(v)\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}1-\cos(m^{\varepsilon}_{*}+\sqrt{\varepsilon}v)dx\Big\}$
@@
$\hskip 76.82243pt\cdot\exp\Big\{-\frac{1}{2\varepsilon}\|\partial_{x}m^{\varepsilon}_{*}\|_{L^{2}}^{2}-\frac{1}{\varepsilon}\langle(-\partial_{x}^{2})m^{\varepsilon}_{*},\sqrt{\varepsilon}v\rangle_{L^{2}}\Big\}$
@@
(7.2) $\hskip 76.82243pt\cdot\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)\,\mu^{\perp}_{\xi_{1},\dots,\xi_{Q}}(dv)\,d\sigma(\xi_{1},\dots,\xi_{Q}),$
@@
where $\mu^{\perp}_{\xi_{1},\dots,\xi_{k}}$ is the Brownian bridge with covariance $(-\partial_{x}^{2})^{-1}$ from 0 to 0, on the normal space, and $d\sigma$ is the surface measure defined in (4.19).
Note that the terms in the exponential together with the Cameron-Martin term in $\mu^{\perp}_{\xi_{1},\dots,\xi_{k}}(dv)$ is just $\frac{1}{\varepsilon}E_{L_{\varepsilon}}(\phi)$ where $E_{L_{\varepsilon}}(\phi)=\frac{1}{2}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}\phi|^{2}dx+\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)\,dx$. Taylor expansion gives
$-\frac{1}{\varepsilon}E_{L_{\varepsilon}}(m^{\varepsilon}_{*}+\sqrt{\varepsilon}v)$ $=-\frac{1}{\varepsilon}E_{L_{\varepsilon}}(m^{\varepsilon}_{*})-\frac{1}{\varepsilon}\langle\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{*}),\sqrt{\varepsilon}v\rangle$
@@
(7.3) $\hphantom{X}-\frac{1}{2\varepsilon}\langle\sqrt{\varepsilon}v,(-\partial_{x}^{2}+\cos m^{\varepsilon}_{*})\sqrt{\varepsilon}v\rangle+\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v),$
where $\mathcal{E}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)=\int_{-L_{\varepsilon}}^{L_{\varepsilon}}O\big((\sqrt{\varepsilon}v)^{3}\big)dx$. Therefore
$Z[F]$ $=\int\cdots\int_{U_{\varepsilon}}F(v)\exp\Big\{-\frac{1}{\varepsilon}E_{L_{\varepsilon}}(m^{\varepsilon}_{*})-\frac{1}{\varepsilon}\langle\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{*}),\sqrt{\varepsilon}v\rangle\Big\}$
@@
$\hphantom{XXXXX}\cdot\exp\Big\{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)\Big\}\exp\Big\{-\frac{1}{2}\langle v,\cos(m^{\varepsilon}_{*})v\rangle\Big\}$
(7.4) $\hphantom{XXXXX}\cdot\textup{Det}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)\,\mu^{\perp}_{\xi_{1},\dots,\xi_{Q}}(dv)\,d\sigma(\xi_{1},\dots,\xi_{Q}).$
We estimate the terms in (7.4) one by one, and all the estimates will be uniform in $\xi_{1},\dots\xi_{Q}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}=\big|\log(\varepsilon\log\frac{1}{\varepsilon})\big|$.
By Lemma 3.4 and (4.6),
$\exp\Big\{-\frac{1}{\varepsilon}E_{L_{\varepsilon}}(m^{\varepsilon}_{*})\Big\}$ $=\exp\Big\{-\frac{1}{\varepsilon}|Q|E_{\textup{kink}}-\frac{1}{\varepsilon}O(e^{-cd_{\varepsilon}})\Big\}$
(7.5) $=\exp\Big\{-\frac{1}{\varepsilon}|Q|E_{\textup{kink}}(1+O(\varepsilon^{0+}))\Big\},$
By Lemma 3.6 and the condition $\|\sqrt{\varepsilon}v\|_{L^{2}}\leq\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\tfrac{1}{\varepsilon}}$,
(7.6) $\bigg|\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{*})\sqrt{\varepsilon}vdx\bigg|\leq\frac{1}{\varepsilon}\|\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{*})\|_{L^{2}}\|\sqrt{\varepsilon}v\|_{L^{2}}\lesssim\frac{1}{\varepsilon}e^{-(1-\eta)d_{\varepsilon}}\delta_{\varepsilon}=O(\varepsilon^{\frac{1}{2}-}),$
which implies that
(7.7) $\exp\Big\{-\frac{1}{\varepsilon}\langle\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{*}),\sqrt{\varepsilon}v\rangle\Big\}=1+O(\varepsilon^{\frac{1}{2}-}).$
We now study the second-order term $\frac{1}{2}\langle v,(-\partial_{x}^{2}+\cos m^{\varepsilon}_{*})v\rangle$ in (7.3), which defines a new base Gaussian measure on the normal space. Note that as observed in (5.1)
$-\partial_{x}^{2}+\cos m^{\varepsilon}_{*}=-\partial_{x}^{2}+1-2\sum_{j=1}^{Q}\text{sech}^{2}(\cdot-\xi_{j})+O(e^{-cd_{\varepsilon}})=\mathcal{L}_{\xi_{1},\dots,\xi_{Q}}+O(e^{-cd_{\varepsilon}}),$
where $\mathcal{L}_{\xi_{1},\dots,\xi_{Q}}$ is defined in (5.5). Since $v\in V_{\xi_{1},\dots,\xi_{Q}}$ satisfies the orthogonality conditions, from Lemma 5.4, we can define the Gaussian measure
(7.8) $d\nu^{\perp}_{\xi_{1},\dots,\xi_{Q}}(v)=Z_{\xi_{1},\dots,\xi_{Q}}^{-1}\exp\Big\{-\frac{1}{2}\langle v,\mathcal{L}_{\xi_{1},\dots,\xi_{Q}}v\rangle\Big\}\prod_{x\in[-L_{\varepsilon},L_{\varepsilon}]}dv(x),$
Finally, for the surface measure $d\sigma(\xi_{1},\dots,\xi_{Q})$ on the multi-soliton manifold, since the tangent vectors $\{\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}\}_{j=1}^{Q}$ are almost orthogonal, as in (4.20) the determinant of Jacobian equals
(7.9) $Q\|\partial_{x}m\|_{L^{2}(\mathbb{R})}^{2}(1+O(e^{-cd_{\varepsilon}})).$
By using (7.2), (7.4), (7.5), (7.7), (7.8), and (7.9), we can take out terms independent of tangential modes $\xi_{1},\dots,\xi_{Q}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}$ as follows:
$Z[F]$ $=J_{\varepsilon}\int\cdots\int_{U_{\varepsilon}}F(v)e^{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)}\textup{Det}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)\,d\nu^{\perp}_{m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}}(v)\,d\xi_{Q}\dots d\xi_{1},$